# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v2

@@ 5.2. RQ2 Findings
$\Rightarrow$ Implication 9.
User reviews highlight key features that developers should preserve when updating their apps. Prior work provided approaches that facilitate competitor analysis (Assi et al., 2025a; Hassan et al., 2022; Assi et al., 2021b). For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
@@
These findings highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind. Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
@@
Summary of RQ 2 From our analysis of user reviews across diverse Gen-AI app categories, we find that users are increasingly experimenting with prompts, offering feedback, and forming emotional connections with non-human agents. They expect fast, creative, and context-aware responses, along with personalization and transparency. For developers, this means moving beyond baseline functionality to support user control (e.g., ability to disable Gen-AI features) and differentiated value compared with AI tools (e.g., ChatGPT).
@@ 6.1. Experiment Setup
    Features extraction. Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash and Koot, 2009). Feature selection enhances interpretability, aligns with research goals  (Kreienkamp et al., 2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable). For each topic category $(T)$, we compute the slope $m$ and the standard deviation $\sigma$ for both $Avg_{(T,P)}$ and $Pr_{(T,P)}$. Selecting these features supports more meaningful clustering outcomes by capturing both the directional trend $m$ and variability $\sigma$ of each topic’s trend. It further enables the grouping of topics that exhibit similar temporal behaviours, such as consistently increasing or decreasing and fluctuating or stable patterns.
@@
    Clustering algorithm. K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren Liao, 2005). We apply K-Means clustering to group topics based on their slope $m$ and standard deviation $\sigma$ values. We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun et al., 2023; Assi et al., 2025b). K-Means is suitable for low-dimensional feature spaces (Ikotun et al., 2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
@@
    Optimal number of clusters. To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari et al., 2023), across multiple values of $k$. Based on the silhouette results, we select the $k$ value closest to 1. The optimal silhouette scores correspond to $k$=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$. We complement this with qualitative validation through manual inspection, confirming that $k$=3 yields the most interpretable and distinct groupings.
Step 3: manual trend investigation. To interpret the identified clusters, we perform qualitative analysis on random review samples from key time periods. For each selected topic category, we identify three representative intervals, typically from the beginning, a period of fluctuation, and the end of the timeline. From each interval, we extract a statistically representative random sample of reviews using a 95% confidence level and a 10% margin of error. We manually code the sampled reviews to identify recurring user concerns, behavioural patterns, and shifts in expectations. This process provides qualitative context that enhances our understanding of the observed quantitative trends.
@@ 6.2. Findings
Observation 3.1: Gen-AI topics consistently receive higher user ratings than non-Gen-AI topics, with an average difference of 0.8 rating stars over time. Specifically, the average rating for Gen-AI topics is 4.0, compared to 3.2 for non-Gen-AI topics. As illustrated in Figure 18, the percentage of reviews discussing Gen-AI topics increases over time. The percentage of Gen-AI reviews increases by 25%, from 18% in 2022-H1 to 43% in 2024-H2, reflecting a growing user focus on Gen-AI features. Despite this surge in volume, the average rating for Gen-AI topics remains stable over time, as shown in Figure 19 with only a slight change of ± 0.2 stars across the observed period. Upon further investigation, we find that this stability is the result of diverging trends across individual topic categories. As illustrated in Figure 20, some topic categories (e.g., Content Policy & Censorship) exhibit increasing trends, others (e.g., Content Quality) decline, while several topics (e.g., AI Performance) remain stable. These opposing directional trends counterbalance each other when aggregated, leading to the observed overall stability in Gen-AI topic ratings.
Figure 18. Evolution of the Percentage of Gen-AI Reviews
Figure 19. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics
Gen-AI topics analysis. As established in our experimental setup, $k=3$ yields the most interpretable and well-separated clusters. These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 20 and 21 for the average ratings and the percentage of reviews, respectively.
@@
Figure 20. Trend Clusters for Average Ratings of Gen-AI Topics
Figure 21. Trend Clusters for Percentage of Gen-AI Topics
@@
We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar, 2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity. Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino and Gudmundsson, 2024). Hence, the decline in the average rating of Content Quality can be explained by the change in observed characteristics between early adopters and mainstream users.
$\Rightarrow$ Implication 12.
Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users. Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust.
Observation 3.5: The AI Performance topic exhibits a stable trend in average user ratings over time, characterized by a slight increase in 2023-H1, and followed by a modest decline in 2024-H1. To account for this overall stability, we analyze fluctuations at the app category level. We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others. This inter-category balancing effect contributes to the relatively flat trend observed in the aggregate ratings.
@@
$\Rightarrow$ Implication 13.
AI integration should be purpose-driven and context-aware. Raji et al. (Raji et al., 2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility. The authors argue that the assumption of AI’s inherent value can lead to ineffective or even harmful implementations, stressing the need for critical assessment before integration. Our findings further compel developers to critically assess whether Gen-AI meaningfully enhances a given task, rather than assuming its value by default. Inappropriate or unnecessary use of AI can degrade usability and frustrate users.
Observation 3.9: The average user ratings for the Content Policy & Censorship topic follow an increasing trend, with fluctuations in 2022-H2 and 2024-H1. We observe that user ratings decrease in 2022-H2, rise steadily through 2023-H2, and decrease slightly in 2024-H1. To understand this trend, we examine the raised user concerns in three random samples of user reviews in 2022-H2, 2023-H2, and 2024-H1. We find that content policy complaints discuss three main concerns: 1) ethical AI training practices, 2) the effectiveness of content filters, and 3) false-positive censorship.

### v3

@@ 5.2. RQ2 Findings
$\Rightarrow$ Implication 4.
ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers. Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search. The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary). User reviews highlight key features that developers should preserve when updating their apps. Prior work provided approaches that facilitate competitor analysis (Maram_LLM_Cure; PeerApps_SANER; Assi_Hassan_Tian_Zou_2021). For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
@@
The findings further highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind. Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
Observation 2.4: Developer Responses to Gen-AI Reviews Are Primarily Courteous or Diagnostic, With Few Immediate Resolutions. Table 8 shows that for Gen-AI apps, the most common developer response type is Acknowledgment and Appreciation (31%), which consists largely of brief thank-you messages. These replies most often accompany positive reviews and serve as courteous engagement rather than substantive technical follow-up.
The second most frequent category, Request for Additional Information (30%), reflects the practical limits of troubleshooting Gen-AI issues through review threads. Even when space is not a constraint, users rarely provide the detailed context needed to diagnose problems that relate to model behavior, device configuration, or account-specific issues. Developers therefore redirect users to email or in-app support channels where logs, device information, or credentials can be shared securely.
By contrast, Issue Resolved and Update Advised appears rarely (2%), whereas Acknowledged and Under Investigation is more common (10%). This pattern aligns with the nature of Gen-AI systems: many issues require model adjustments, infrastructure changes, or updates to safety and moderation mechanisms, none of which can be addressed immediately or within the constraints of a review thread. As a result, developers are more likely to acknowledge problems and communicate that work is underway than to confirm the issue has been resolved.
Table 8. Summary of developer response types, including descriptions, number of reviews, and percentage distribution (Representative subset considered).
Topic Description # Reviews %
Acknowledgment and Appreciation Generic responses that thank users for their feedback, often including a brief note about considering future updates or improvements. 30 31%
Request for Additional Information Responses indicating that the reported issue lacks sufficient detail and requesting that the user contact the developer directly for further clarification or resolution. 29 30%
Clarification or Policy Rebuttal Responses that clarify or refute user concerns, often explaining app policies or correcting misunderstandings. Most common for reviews concerning free versus premium features. 13 14%
Guidance and Troubleshooting Tips Responses that address reported issues by offering troubleshooting advice, resources, or specific steps to help users resolve the problem. 12 13%
Acknowledged and Under Investigation Responses confirming that the development team is aware of the issue and is currently investigating or working on a solution. 10 10%
Issue Resolved and Update Advised Responses informing the user that the reported issue has been resolved and recommending that they update or reinstall the app. 2 2%
Total 96 100%
Hassan et al. (developer_response_safwat) analyzed developer replies across a broad set of mobile apps and identified recurring response drivers such as “thank the user”, “ask for more details”, “provide guidance”, “ask for endorsement”, “advertise other products”, wait for updates, and “justify advertisements”. Many of these practices align closely with our taxonomy. For example, their “thank the user” category maps to our “Acknowledgment and Appreciation,” their “ask for more details” category maps to our “Request for Additional Information,” and their “justify the advertisements” category maps onto our broader Clarification or Policy Rebuttal, which includes explanations about subscription models, premium features, and content restrictions. Notably, their work does not treat “issue resolved” as a dominant response driver, and we similarly observe that confirmed resolutions are relatively rare in our Gen-AI dataset (2%). We report the category explicitly for completeness of Gen-AI developer behaviors, but its low prevalence mirrors prior findings rather than contrasting with them. Overall, the strong correspondence between the two taxonomies suggests that general response practices extend to Gen-AI apps, while our domain-specific categories capture the additional emphasis developers place on clarifying AI-related behavior and managing expectations around the evolving nature of Gen-AI features.
Observation 2.5: App categories show highly distinctive developer response patterns. Table 9 reveals variation in how developers across different app categories respond to user feedback. Some categories exhibit exceptionally high engagement, while others show very limited interaction. For example, Games stands out with both the highest reply ratio (56%) and the longest average response delay (44.7 days). In contrast, Education demonstrates the shortest response times, with developers replying in less than a day on average (0.7 days). Meanwhile, Entertainment apps show the lowest reply ratio among categories with meaningful activity (2%). These contrasting patterns suggest that developer responsiveness is shaped not only by category size but also by the operational demands, user expectations, and support norms unique to each domain.
Table 9. Developer response activity across app categories, including total reviews, total developer replies, overall reply ratio, and average response delay (Full dataset considered).
Category Total Reviews Total Replies Reply Ratio Avg Delay (days)
Productivity 169,119 8,407 5% 8.4
Photography 126,277 31,070 25% 4.5
Art & Design 125,362 39,593 32% 5.5
Entertainment 118,669 2,783 2% 1.7
Education 41,905 16,034 38% 0.7
Video Players & Editors 36,507 11,123 30% 2.2
Tools 35,046 905 3% 32.45
Music & Audio 9,150 2,127 23% 3.4
Health & Fitness 7,849 338 4% 7.7
Games 5,070 2,848 56% 44.7
Books & Reference 1,112 1 0% 1.0
All Categories 676,066 115,164 20% 10.2
Our manual analysis of developer responses in the Games category helps explain why it exhibits both the highest reply ratio and the longest response delay. Responses in which developers report that an issue has been fixed account for 30% of all replies in this category, far exceeding the 2% observed in the overall dataset. This emphasis on confirmed fixes aligns with the long response delay shown in Table 9. This likely reflects that developers often wait until a patch is deployed before responding, resulting in slower but more substantive communication. Responses indicating that an issue is acknowledged and under investigation represent 13% of replies, slightly higher than the overall 10%. Together, these categories show that Games developers place strong emphasis on technical issue resolution and ongoing investigation.
Generic thank-you messages appear in only 12% of replies, considerably lower than the overall 31%. The remaining responses highlight active engagement with user issues: 28% request additional information, 10% provide troubleshooting guidance, and 8% offer clarifications or policy rebuttals. Overall, developer communication in Games is markedly problem-oriented and resolution-driven, which leads to both high engagement and extended response timelines.
A closer look at Education apps helps explain why this category exhibits the fastest average developer response time. In our manual sample, the majority of Education responses fall under Acknowledgment and Appreciation (73%), which is markedly higher than both the overall rate of 31% and the 12% observed in Games. These replies are typically short, template-based thank-you messages that require minimal effort to generate, allowing developers to respond quickly and consistently.
Education apps also show a comparatively high rate of Request for Additional Information (16%). This proportion is notably larger than the overall sample and reflects a common pattern in Gen-AI Education reviews. Users frequently report that the app produced incorrect answers, yet they often omit the original prompt, the model’s response, or screenshots of the interaction. Without these details, developers cannot diagnose or reproduce the issue within the review thread. As a result, they quickly reply by directing users to contact support channels where complete context can be shared securely. This response type is therefore both common and fast, since it does not require problem investigation within the review platform itself.
By contrast, response types requiring deeper engagement occur far less frequently in Education, including Guidance and Troubleshooting Tips (2%), Acknowledged and Under Investigation (6%), and Issue Resolved and Update Advised (1%). Together, these patterns show that Education apps tend to rely primarily on quick acknowledgments or requests for additional information, which contributes to the category’s exceptionally short response delay.
Entertainment apps show the lowest reply ratio in our dataset, and the distribution of response types helps explain this pattern. Acknowledgment and Appreciation responses account for only 14%, far below the overall rate of 31%. This indicates that developers in this category do not typically reply to routine or positive reviews and instead reserve responses for situations where user action or developer follow-up is required. Correspondingly, a large share of replies fall under Request for Additional Information (38%) and Guidance and Troubleshooting Tips (22%), suggesting that developers primarily intervene when users report crashes, playback issues, or account problems that require diagnostic detail. The relatively high proportion of Acknowledged and Under Investigation (19%) further shows that many of these reports involve technical issues that cannot be resolved immediately, which may discourage broader engagement with lower-priority feedback. Together, these patterns indicate that Entertainment developers focus their limited response efforts on actionable problem reports rather than general user engagement, resulting in a lower overall reply ratio.
Observation 2.6: Developer responsiveness varies substantially across Gen-AI topic categories. Table 10 summarizes developer reply behavior for each Gen-AI topic category. The topic with the highest reply ratio is Accessibility & Inclusivity (47%), despite having the smallest number of reviews. This indicates that developers treat accessibility-related feedback as a high-impact area that warrants immediate and consistent attention. Such reviews often report barriers to use for specific user populations, making them critical from both ethical and product-quality perspectives. The short average delay (0.2 days) further reflects the urgency developers assign to this category.
Beyond accessibility, the next highest reply ratios are found in Utility & Use Cases (18%) and Content Quality (14%), suggesting that developers are highly responsive to feedback about how users apply AI features and whether the generated outputs meet expectations. These topics directly affect user satisfaction and daily usability, which likely motivates more frequent developer engagement.
In contrast, the categories with the lowest reply ratios are Customer Support & Community (1%), Human–AI Interaction & Emotional Connection (2%), and Content Policy & Censorship (5%). These topics often involve complaints about communication quality, emotional tone, safety constraints, or moderation decisions. Such issues are typically systematic, tied to model behavior or policy rules that cannot be quickly adjusted, which may explain the lower response frequency and the longer average delays seen in these categories.
This contrast between highly actionable categories and structurally constrained categories reinforces the idea that responsiveness is shaped by whether an issue is immediately fixable or requires deeper system changes. Overall, these patterns show that developers prioritize issues that are actionable in the short term (e.g., accessibility barriers, inaccurate outputs, task-level failures) while responding less frequently to concerns that require deeper model changes or policy revisions. This reflects a practical triaging strategy driven by the nature of Gen-AI systems and the feasibility of immediate fixes.
Table 10. Developer response patterns across Gen-AI topic categories, including total reviews, total developer replies, reply ratio, and average response delay (Full dataset considered).
Topic Category Total Reviews Total Replies Reply Ratio Avg Delay (days)
AI Performance 61,071 7,152 12% 4.2
Utility & Use Cases 56,505 9,895 18% 0.6
Content Quality 44,594 6,275 14% 4.3
Content Policy & Censorship 27,521 1,397 5% 7.9
Human–AI Interaction & Emotional Connection 25,142 542 2% 7.5
Features & Functionality 16,275 1,892 12% 5.5
Customer Support & Community 3,361 28 1% 10.3
Comparison to Other Apps 2,315 287 12% 6.0
Personalization & Customization 1,283 106 8% 2.1
Accessibility & Inclusivity 96 45 47% 0.2
All Topics 238,168 27,619 12% 4.9
Observation 2.7: Developers respond less frequently to Gen-AI reviews than to non-Gen-AI reviews as illustrated in Figure 18. We compared per-app developer response rates for Gen-AI and non-Gen-AI reviews using a paired non-parametric Wilcoxon signed-rank test ($n=164$ apps). Developers replied to a median of 9.6% of Gen-AI reviews versus 17.7% of non-Gen-AI reviews. Although the median per-app difference is modest (0.7-0.7 percentage points), the effect is statistically significant ($W=2911$, $p<0.001$), indicating a consistent pattern across apps.
This gap aligns with the qualitative patterns observed earlier. Gen-AI reviews often describe failures rooted in model behavior, data quality, or moderation mechanisms, which are difficult to diagnose from short review text. By contrast, non-Gen-AI reviews commonly involve UI issues, feature requests, or account problems that are easier to acknowledge, triage, or resolve within the constraints of the review interface. As a result, developers appear slightly more willing to engage with non-AI feedback where their replies can directly influence user experience. In addition to analyzing user perspectives, we also examine how developers respond to Gen-AI reviews, capturing patterns of engagement, resolution, and support practices across app categories.
Observation 2.8: Developers respond to Gen-AI and non-Gen-AI reviews at similar speeds. Figure 18 compares developer response delays for Gen-AI and non-Gen-AI reviews across all apps. While Gen-AI reviews show a slightly lower median delay (a difference of 0.15-0.15 days), this difference is not statistically significant. In practical terms, once developers choose to reply, they tend to respond at comparable speeds regardless of whether the review discusses Gen-AI functionality or non-AI aspects of the app. This suggests that the lower engagement with Gen-AI reviews observed in Observation 2.7 reflects differences in response likelihood rather than response promptness.
Figure 18. Comparison of developer responsiveness across apps. Left: Relative likelihood of developer response to Gen-AI reviews compared to non–Gen-AI reviews, expressed as the ratio of per-app response probabilities (values below 1 indicate lower responsiveness to Gen-AI reviews). Right: Distribution of mean developer response delays per app for Gen-AI and non–Gen-AI reviews; the two delay distributions do not differ significantly. (Full dataset considered.)
$\Rightarrow$ Implication 6.
Taken together, our developer response findings show that review threads in Gen-AI apps primarily serve as an initial point of contact rather than a channel for resolving AI-specific failures. Developers face structural barriers when responding to Gen-AI feedback, since many issues require information that cannot be shared in public reviews or involve model- or policy-level changes that cannot be deployed quickly. As a result, developers tend to reserve detailed engagement for issues that are actionable in the short term and redirect users elsewhere when deeper diagnostic context is required.
These patterns highlight a broader platform-level limitation: while app stores surface user feedback effectively, they do not provide the depth of interaction required to investigate or clarify Gen-AI behaviours. There is therefore an opportunity for platform owners to introduce more structured, built-in support mechanisms that complement review sections and enable richer, privacy-preserving communication channels. Such improvements would better support both developers and users as Gen-AI apps become more complex and more central to mobile ecosystems.
@@
Summary of RQ 2 From our analysis of user reviews across diverse Gen-AI app categories, we find that users are increasingly experimenting with prompts, offering feedback, and forming emotional connections with non-human agents. They expect fast, creative, and context-aware responses, along with personalization and transparency. For developers, this means moving beyond baseline functionality to support user control (e.g., ability to disable Gen-AI features) and differentiated value compared with AI tools (e.g., ChatGPT). Our developer-response analysis reinforces this point: developers most often engage with actionable, user-reported issues and struggle to address Gen-AI failures directly within the review platform. This suggests that better support channels will be essential for meeting rising user expectations.
@@ 6.1. Experiment Setup
    Features extraction. Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash_Koot_2009). Feature selection enhances interpretability, aligns with research goals  (Kreienkamp_Agostini_Monden_Epstude_DeJonge_Bringmann_2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable). For each topic category $(T)$, we compute the slope $m$ and the standard deviation $\sigma$ for both $Avg_{(T,P)}$ and $Pr_{(T,P)}$. Selecting these features supports more meaningful clustering outcomes by capturing both the directional trend $m$ and variability $\sigma$ of each topic’s trend. It further enables the grouping of topics that exhibit similar temporal behaviours, such as consistently increasing or decreasing and fluctuating or stable patterns.
@@
    Clustering algorithm. K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren_Liao_2005). We apply K-Means clustering to group topics based on their slope $m$ and standard deviation $\sigma$ values. We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023; 10.1145/3721125). K-Means is suitable for low-dimensional feature spaces (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
@@
    Optimal number of clusters. To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari_DwiNugroho_Baraku_NovriYanda_Liwardana_2023), across multiple values of $k$. Based on the silhouette results, we select the $k$ value closest to 1. The optimal silhouette scores correspond to kk=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$. We complement this with qualitative validation through manual inspection, confirming that kk=3 yields the most interpretable and distinct groupings.
Step 3: manual trend investigation. To interpret the identified clusters, we perform qualitative analysis on statistically significant random review samples from key time periods. For each selected topic category, we identify three representative intervals, typically from the beginning, a period of fluctuation, and the end of the timeline. From each interval, we extract a statistically representative random sample of reviews using a 95% confidence level and a 10% margin of error. We manually code the sampled reviews to identify recurring user concerns, behavioural patterns, and shifts in expectations. This process provides qualitative context that enhances our understanding of the observed quantitative trends.
@@ 6.2. Findings
Observation 3.1: Gen-AI topics consistently receive higher user ratings than non-Gen-AI topics, with an average difference of 0.8 rating stars over time. Specifically, the average rating for Gen-AI topics is 4.0, compared to 3.2 for non-Gen-AI topics. As illustrated in Figure 19, the percentage of reviews discussing Gen-AI topics increases over time. The percentage of Gen-AI reviews increases by 25%, from 18% in 2022-H1 to 43% in 2024-H2, reflecting a growing user focus on Gen-AI features.
This growth indicates that users are increasingly evaluating apps through the lens of their Gen-AI features. As AI becomes a central part of the user experience, developers can no longer treat it as an add-on; users expect it to perform reliably and meaningfully enhance the app’s core value. The rising share of Gen-AI reviews therefore signals that user satisfaction and engagement increasingly depend on how well these features are designed, integrated, and maintained.
Despite this surge in volume, the average rating for Gen-AI topics remains stable over time, as shown in Figure 20 with only a slight change of ± 0.2 stars across the observed period. Upon further investigation, we find that this stability is the result of diverging trends across individual topic categories. As illustrated in Figure 21, some topic categories (e.g., Content Policy & Censorship) exhibit increasing trends, others (e.g., Content Quality) decline, while several topics (e.g., AI Performance) remain stable. These opposing directional trends counterbalance each other when aggregated, leading to the observed overall stability in Gen-AI topic ratings. The following observations examine these diverging topic-specific trajectories in greater detail.
Figure 19. Evolution of the Percentage of Gen-AI Reviews (Full dataset considered).
Figure 20. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics (Full dataset considered).
Gen-AI topics analysis. As established in our experimental setup, $k=3$ yields the most interpretable and well-separated clusters. These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 21 and 22 for the average ratings and the percentage of reviews, respectively.
@@
Figure 21. Trend Clusters for Average Ratings of Gen-AI Topics (Full dataset considered).
Figure 22. Trend Clusters for Percentage of Gen-AI Topics (Full dataset considered).
@@
We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar_2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity. Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino2024). Hence, the decline in the average rating of Content Quality can be explained by the change in observed characteristics between early adopters and mainstream users.
Observation 3.5: The AI Performance topic exhibits a stable trend in average user ratings over time, characterized by a slight increase in 2023-H1, and followed by a modest decline in 2024-H1. To account for this overall stability, we analyze fluctuations at the app-category level. We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others, producing the flat global average.
However, this apparent stability conceals divergent user experiences. Categories such as Education and Games show clear improvement, reflecting that users in these domains perceive AI as directly enhancing learning and task performance. In contrast, most other categories (e.g., Productivity, Entertainment, Books & Reference) show gradual declines. This may suggest that developers in these areas are integrating AI features prematurely, before the technology is fully optimized for their use cases. When users accustomed to fast and accurate AI systems encounter slower or less reliable integrations, their satisfaction drops. These declines therefore point to gaps in seamless integration and contextual adaptation rather than outright failure of AI Performance.
For developers, this indicates that AI performance has become a key competitive factor. Users now expect fast, accurate, and context-aware AI behavior. When AI integrations fall short or introduce lag, inconsistency, or irrelevant responses, users quickly express dissatisfaction. Sustaining trust therefore depends on ensuring that AI features operate seamlessly within the app’s core experience rather than as add-ons.
@@
$\Rightarrow$ Implication 7.
Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users. Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust. AI integration should be purpose-driven and context-aware. Raji et al. (Raji_Kumar_Horowitz_Selbst_2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility. The authors argue that the assumption of AI’s inherent value can lead to ineffective or even harmful implementations, stressing the need for critical assessment before integration. Our findings further compel developers to critically assess whether Gen-AI meaningfully enhances a given task, rather than assuming its value by default. Inappropriate or unnecessary use of AI can degrade usability and frustrate users.
Observation 3.9: The average user ratings for the Content Policy & Censorship topic follow an increasing trend, with fluctuations in 2022-H2 and 2024-H1.
To understand this trend, we examine the raised user concerns in three random samples of user reviews in 2022-H2, 2023-H2, and 2024-H1. We find that content policy complaints discuss three main concerns: 1) ethical AI training practices, 2) the effectiveness of content filters, and 3) false-positive censorship.
This pattern reflects how content policy concerns were especially prominent in the early stages of Gen-AI deployment, when moderation systems were still limited and users frequently encountered unsafe or poorly filtered outputs. Over time, developers appear to have become increasingly aware of these issues and implemented more effective safeguards, leading to gradual improvement in user satisfaction. The upward trajectory in ratings thus signals a maturing ecosystem where content safety and moderation mechanisms are continuously refined. For developers, this emphasizes the importance of iterative moderation design. Continuous refinement of content filters not only improves safety but also builds long-term user trust.